OpenStudy (anonymous):

Two spherical pieces of cookie dough have radii of 3cm & 5cm. The pieces are combined to form one large spherical piece of dough. What is the approximate radius of the new sphere of dough?

OpenStudy (anonymous):

The traffic cone is 19 inches tall & has a radius of 5 inches. Find the lateral area.

OpenStudy (anonymous):

You'll need to find the volume of the first two balls of dough. The combined ball will have the sum of the two volumes. \[V_1=\frac{4}{3}\pi (3)^3=36\pi~~~~\text{and}~~~~V_2=\frac{4}{3}\pi (5)^3=\frac{500}{3}\pi\] so the third ball would have volume \(V_3=\dfrac{608}{3}\pi\). Solve for the third ball's radius: \[\frac{608}{3}\pi=\frac{4}{3}\pi r^3~~\iff~~152=r^3~~\iff~~r\approx5.337\]

OpenStudy (anonymous):

The traffic cone is 19 inches tall & has a radius of 5 inches. Find the lateral area.